Solve the following system of linear equations. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)
x+y=6
2x+2y=-2
x = ?
y = ?
Pls show me the steps on how to get the answer. thanks

Answer by stanbon(75887)

x+y=6
2x+2y=-2
-------------
Divide thru 2nd equation by 2 to get:
x+y = 6
x+y = -1
-----------
Subtract to get:
0 = 7
----
That is a contradiction which resulted from
assuming the x's and the y's of the two equations
have the same value.
They do not.
There is no point (x,y) which satisfies both equation.
The lines are parallel.
---
Solution: NONE
======================================
Cheers,
Stan H.
=================



x = ?
y = ?